The Campbell Company is considering adding a robotic paint sprayer to its production line. The sprayer's base price is $1,170,000, and it would cost another $21,500 to install it. The machine falls into the MACRS 3-year class (the applicable MACRS depreciation rates are 33.33%, 44.45%, 14.81%, and 7.41%), and it would be sold after 3 years for $579,000. The machine would require an increase in net working capital (inventory) of $10,500. The sprayer would not change revenues, but it is expected to save the firm $498,000 per year in before-tax operating costs, mainly labor. Campbell's marginal tax rate is 30%.

What is the Year 0 net cash flow?
$



What are the net operating cash flows in Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar.

Year 1 $
Year 2 $
Year 3 $

What is the additional Year 3 cash flow (i.e, the after-tax salvage and the return of working capital)? Do not round intermediate calculations. Round your answer to the nearest dollar.
$

If the project's cost of capital is 11 %, what is the NPV of the project? Do not round intermediate calculations. Round your answer to the nearest dollar.
$

Should the machine be purchased?
Answer #1
NET OPERATING CASH FLOW
Year 0 1 2 3
a Base Price $1,170,000
b Installation Cost $21,500
c=a+b Total Asset Cost $1,191,500
d Net working capital cash out flow $10,500
e=c+d Net Cash Outflow $1,202,000
A Depreciation Rate 33.33% 44.45% 14.81% 7.41%
B=1191500*A Annual Depreciation $397,127 $529,622 $176,461
C Accumulated Depreciation $397,127 $926,749 $1,103,210
D=1191500-C Book value at end of year $794,373 $264,751 $88,290
E=B*30% Depreciation Tax Shield $119,138 $158,887 $52,938
F Before tax annual savings $498,000 $498,000 $498,000
G=F*(1-0.3) After tax annual savings $348,600 $348,600 $348,600
H=E+G Operating Cash Flow $467,738 $507,487 $401,538
Additional year 3 cash flow:
I Release of net working capital $10,500
J Salvage value of equipment $579,000
K Book value of equipment $88,290
L=J-K Gain on sale $490,710
M=L*0.3 Tax on Gain $147,213
P=J-M After tax Cash flow on salvage $431,787
Q=I+P Additional year 3 cash flow: $442,287
Year 0 net Cash Flow ($1,202,000)
Net Operating cash Flow
Year1 $467,738
Year2 $507,487
Year3 $401,538
Additional Year 3 Cash Flow $442,287
Total Cash Flow in year 3 $843,825 (401538+442287)
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=11%=0.11
N=Year of Cash Flow
N Year 0 1 2 3
R Cash Flow ($1,202,000) $467,738 $507,487 $843,825 SUM
PV=R/(1.11^N) Present Value (PV) of Cash Flow: ($1,202,000) $421,386 $411,887 $616,998 $248,271
NPV of the Project $248,271
YES, THE MACHINE SHOULD BE PURCHASED
NPV is Positive